The sum of three numbers is 123. The second number is 9 less than two times the first number. The third number is 6 more than three times the first number. Find the three numbers.
Answer by pwac(253)

Let x= first number
2x-9=second number
3x+6=third number so if we add these terms together
x+2x-9+3x+6=123
6x-3=123 add 3 to both sides
6x=126 divide both sides by 6
x=21
2x-9=33
3x+6=69
21+33+69=123 correct
Pete